# Two Characterizations of Geometrically Infinite Actions on Gromov Hyperbolic Spaces

### v1

@@ 4. Uncountably Many Non-conical Points
Before turning to the proof, we note that the strategy originating in [BIS96] (and further developed in [KL19]) has by now become standard. However, in the level of generality considered here, a few additional arguments are needed. In particular, one must verify that the geodesic rays produced by this method are genuinely escaping, so that their endpoints are non-conical.
@@
Throughout this section, let $G$ be a discrete group that admits a non-elementary, geometrically infinite, proper action on a proper $\delta$-hyperbolic space $X$ for some $\delta\geq 0$, and fix a basepoint $o\in X$. By Lemma 2.3, there exist constants $L>16\delta$, $\lambda\geq 1$, and $C\geq 0$ such that every $L$-local $(3,2D_{0})$-quasi-geodesic in $X$ is a $(\lambda,C)$-quasi-geodesic. By Lemma 2.7, there exists a constant $B>0$ such that $d_{\text{Haus}}(p,[p_{-},p_{+}])\leq B$ for every $(\lambda,C)$-quasi-geodesic $p$.
@@ 4.1. Existence of escaping hyperbolic elements in one direction
Figure 3.
@@
The following result plays the role of [KL19, Lemma 4.1].
@@
Since $\lim\limits_{n\to\infty}d(\gamma_{n},Go)=\infty$, and since $x_{n},p_{n}]\subset\gamma_{n}$ and $p_{n+1},y_{n+1}]\subset\gamma_{n+1}$, the terms $d(Go,[x_{n},p_{n}])$ and $d(Go,[p_{n+1},y_{n+1}])$ tend to infinity.
@@ 4.2. Construction of escaping geodesic rays
It remains to show that $\gamma_{+}\in\Lambda^{nc}G$. Note that the Hausdorff distance between $p_{n}$ and a subpath of a translate of $\gamma_{n}$ is at most $D_{0}$. Since $d(\gamma_{n},Go)\to\infty$, it follows that $d(p_{n},Go)\to\infty$. Because $q_{n}$ is a translate of $x_{n},y_{n+1}]$, Lemma 4.2 implies $d(q_{n},Go)\to\infty$. Therefore, by the definition of conical points, $\gamma_{+}$ is non-conical. ∎
@@
$d_{\text{Haus}}(p([\lambda(2D_{0}+200\delta+B)+C,\infty)),[y,\gamma_{+}])\leq B.$

### v2

@@ 4. Uncountably Many Non-conical Points
Before turning to the proof, we note that the strategy originating in [BIS96] (and further developed in [KL19]) has by now become standard. However, at the level of generality considered here, a few additional arguments are needed. In particular, one must verify that the geodesic rays produced by this method are genuinely escaping, so that their endpoints are non-conical.
@@
Throughout this section, let $G$ be a discrete group that admits a non-elementary, geometrically infinite, proper action on a proper $\delta$-hyperbolic space $X$ for some $\delta\geq 0$, and fix a basepoint $o\in X$. By Lemma 2.13, there exist constants $L>16\delta$, $\lambda\geq 1$, and $C\geq 0$ such that every $L$-local $(3,2D_{0})$-quasi-geodesic in $X$ is a $(\lambda,C)$-quasi-geodesic. By Lemma 2.12, there exists a constant $B>0$ such that $d_{\mathrm{Haus}}(p,[p_{-},p_{+}])\leq B$ for every $(\lambda,C)$-quasi-geodesic $p$.
@@ 4.1. Existence of escaping hyperbolic elements in one direction
Figure 3. Choosing the escasping sequence of axes.
@@
The following result corresponds to [KL19, Lemma 4.1].
@@
Since $\lim\limits_{n\to\infty}d(\gamma_{n},Go)=\infty$, and since $x_{n},p_{n}]\subset\gamma_{n}$ and $p_{n+1},y_{n+1}]\subset\gamma_{n+1}$, the terms $d(Go,[x_{n},p_{n}])$ and $d(Go,[p_{n+1},y_{n+1}])$ tend to infinity. See Figure 4.
@@ 4.2. Construction of escaping geodesic rays
It remains to show that $\gamma_{+}$ is not conical. Note that the Hausdorff distance between $p_{n}$ and a subpath of a translate of $\gamma_{n}$ is at most $D_{0}$. Since $d(\gamma_{n},Go)\to\infty$, it follows that $d(p_{n},Go)\to\infty$. Because $q_{n}$ is a translate of $x_{n},y_{n+1}]$, Lemma 4.2 implies $d(q_{n},Go)\to\infty$. Therefore, by the definition of conical points, $\gamma_{+}$ is non-conical. ∎
@@
$d_{\mathrm{Haus}}(p([\lambda(2D_{0}+200\delta+B)+C,\infty)),[y,\gamma_{+}])\leq B.$